you are playing a card game called crime scene. thare are 3 diamond cards, 9handcuffs card and 5 car cards. You can draw 2 cards without replace ment. if you draw one diamond card and one car card, you win the game. For the following questions round your answer to 2 decimal places.
Find the probability that your first card is a diamond. Find the probability that your first card is not a diamond.
Find the probability that both your cards are handcuffs.
Given that you draw a diamond first find the probability that you win the game.
Answer #1

1)

probability that your first card is a diamond =3/17 =0.18

2)

probability that your first card is not a diamond.=1-3/17 =1-0.18 =0.82

3)

probability that both your cards are handcuffs=(9/17)*(8/16)=0.26

4)

P(win |first card was diamond )=P(2nd card is car card) =5/16 =0.31